# D³-Predictor: Noise-Free Deterministic Diffusion for Dense Prediction

### v2

@@ 2.1 Diffusion-based Dense Prediction
Several works have explored leveraging diffusion models for dense prediction. Methods such as DIFT [tang2023emergent] and CleanDIFT [stracke2025cleandift] extract diffusion priors as general-purpose visual features and train lightweight probes for dense prediction tasks, but they are not end-to-end and their performance remains limited. In contrast, other methods [ke2024repurposing, gui2025depthfm, zhao2025diception, fu2024geowizard] fine-tune a pretrained diffusion model for dense prediction tasks, achieving impressive results. However, these fine-tuning methods rely on stochastic noise to exploit diffusion priors. GenPercept [xu2024matters] and E2E-FT [garcia2025fine] attempt to alleviate stochasticity but disrupt the integrity of diffusion priors, while StableNormal [ye2024stablenormal] suppresses stochasticity at the cost of additional computational overhead. In contrast, our method reformulates the diffusion model into a noise-free framework to better suit dense prediction in a simple self-supervised manner, ensuring both prediction performance and speed.
@@ 3.2.1 Self-Supervised Diffusion Priors Aggregation
The core of our approach is to train our D³-Predictor $f_{D^{3}}(\cdot)$ to incorporate a noise-free and complete diffusion prior compared to the pre-trained diffusion model, which is enhanced to be more suitable for deterministic dense prediction tasks. Inspired by the noise-free diffusion feature extraction approach proposed by CleanDIFT[stracke2025cleandift], we achieve this enhancement through a simple self-supervised manner, and crucially, relying solely on the model itself rather than costly external guidance (e.g., DINO [oquab2023dinov2]) adapted by previous works [ye2024stablenormal]. We begin by initializing D³-Predictor with a pretrained diffusion model.
As shown in Fig 3 (a), the insight here is to view the diffusion model as an ensemble of $t$ visual experts, where each expert $f_{\text{exp}}(\cdot,\epsilon,t)$ exhibits a distinct timestep-dependent diffusion prior. While the visual experts take a noisy image $(\epsilon,x)$ together with its timestep $t$ as input, our D³-Predictor operates directly on a clean image $x$. During training, D³-Predictor aligns its internal representations $r_{D^{3}}^{K}$ = $f_{D^{3}}^{K}(x)$ with those of $t$ visual experts $r_{\text{exp}}^{K,t}$ = $f_{\text{exp}}^{K}(x,\epsilon,t)$. Here, $K$ is the index set of model layers selected for alignment.
To stabilize this timestep-dependent representation alignment, we introduce a lightweight, trainable timestep-conditioned projection head $P_{\varphi}^{K}(\cdot,t)$ parameterized by $\varphi$, as shown in Fig 3 (a). This projection head serves to project the single, noise-free representation $r_{D^{3}}^{K}$ into $T$ distinct timestep-dependent representation spaces for the purpose of matching the expert representations. This mechanism implicitly requires our model’s representation to encode the diverse diffusion priors corresponding to different noise levels. We measure the alignment by their point-wise distance:

### v3

@@ 2.1 Diffusion-based Dense Prediction
Several works have explored leveraging diffusion models for dense prediction. Methods such as DIFT [tang2023emergent] and CleanDIFT [stracke2025cleandift] extract diffusion priors as general-purpose visual features and train lightweight probes for dense prediction tasks, but they are not end-to-end and their performance remains limited. While CleanDIFT [stracke2025cleandift] treats the diffusion model as a series of feature extractors, our method views it as an ensemble of visual experts directly optimized for dense prediction via end-to-end, task-specific supervision. On the other hand, other methods [ke2024repurposing, gui2025depthfm, zhao2025diception, fu2024geowizard] fine-tune a pretrained diffusion model for dense prediction tasks, achieving impressive results. However, these fine-tuning methods rely on stochastic noise to exploit diffusion priors. GenPercept [xu2024matters] and E2E-FT [garcia2025fine] attempt to alleviate stochasticity but disrupt the integrity of diffusion priors, while StableNormal [ye2024stablenormal] suppresses stochasticity at the cost of additional computational overhead. In contrast, our method reformulates the diffusion model into a noise-free framework to better suit dense prediction in a simple self-supervised manner, ensuring both prediction performance and speed.
@@ 3.2.1 Self-Supervised Diffusion Priors Aggregation
The core of our approach is to train our D³-Predictor $f_{D^{3}}(\cdot)$ to incorporate a noise-free and complete diffusion prior compared to the pre-trained diffusion model, which is enhanced to be more suitable for deterministic dense prediction tasks. We achieve this enhancement through a simple self-supervised manner, and crucially, relying solely on the model itself rather than costly external guidance (e.g., DINO [oquab2023dinov2]) adapted by previous works [ye2024stablenormal]. We begin by initializing D³-Predictor with a pretrained diffusion model.
As shown in Fig 3 (a), the insight here is to view the diffusion model as an ensemble of $t$ visual experts, where each expert $f_{\text{exp}}(\cdot,\epsilon,t)$ exhibits a distinct timestep-dependent diffusion prior. We inherit the mechanism of aligning a clean backbone with a frozen diffusion model from CleanDIFT [stracke2025cleandift]. While the visual experts take a noisy image $(\epsilon,x)$ together with its timestep $t$ as input, our D³-Predictor operates directly on a clean image $x$. During training, D³-Predictor aligns its internal representations $r_{D^{3}}^{K}$ = $f_{D^{3}}^{K}(x)$ with those of $t$ visual experts $r_{\text{exp}}^{K,t}$ = $f_{\text{exp}}^{K}(x,\epsilon,t)$. Here, $K$ is the index set of model layers selected for alignment.
To stabilize this timestep-dependent representation alignment, we introduce a lightweight, trainable timestep-conditioned projection head $P_{\varphi}^{K}(\cdot,t)$ parameterized by $\varphi$. While the use of projection heads for representation alignment has proven effective in prior works [yu2024representation, wu2025representation, jiang2025no, stracke2025cleandift], we specifically leverage the timestep-conditioned projection head adopted by CleanDIFT [stracke2025cleandift] and shown in Fig 3 (a). This projection head serves to project the single, noise-free representation $r_{D^{3}}^{K}$ into $T$ distinct timestep-dependent representation spaces for the purpose of matching the expert representations. This mechanism implicitly requires our model’s representation to encode the diverse diffusion priors corresponding to different noise levels. We measure the alignment by their point-wise distance:
@@ 5 Conclusion
Note: An anonymous static site with our visualization results is available at ./index.html. We will keep updating it and release it publicly with our resources.